# TTT: A Temporal Refinement Heuristic for Tenuously Tractable Discrete Time Reachability Problems

### v1

@@ Abstract
Reachable set computation is an important tool for analyzing control systems. Simulating a control system can show that the system is generally functioning as desired, but a formal tool like reachability analysis can provide a guarantee of correctness. For linear systems, reachability analysis is straightforward and fast, but as more complex components are added to the control system such as nonlinear dynamics or a neural network controller, reachability analysis may slow down or become overly conservative. To address these challenges, much literature has focused on spatial refinement, e.g., tuning the discretization of the input sets and intermediate reachable sets. However, this paper addresses a different dimension: temporal refinement. The basic idea of temporal refinement is to automatically choose when along the horizon of the reachability problem to execute slow symbolic queries which incur less approximation error versus fast concrete queries which incur more approximation error. Temporal refinement can be combined with other refinement approaches and offers an additional “tuning knob” with which to trade off tractability and tightness in approximate reachable set computation. Here, we introduce an automatic framework for performing temporal refinement and we demonstrate the effectiveness of this technique on computing approximate reachable sets for nonlinear systems with neural network control policies. We demonstrate the calculation of reachable sets of varying approximation error under varying computational budget and show that our algorithm is able to generate approximate reachable sets with a similar amount of error to the baseline approach in 20-70% less time.
{IEEEkeywords}
3-5 keywords
1 Introduction
@@
Every controller needs to be analyzed for performance and reliability. Simulation is useful to get an idea of the general behavior of a control policy, but only formal analysis can guarantee correctness with respect to a model of the system and a specification. In this paper, we focus on reach-avoid specifications which dictate a goal set the system must reach and/or an avoid set it must not reach. For a given discrete time dynamical system with a control policy and set of possible starting states the forward reachable set at time $t$ is the set of states that the system could reach $t$ steps into the future. Given a final set instead of an initial set, the $t$-step backward reachable set is the set of states that will reach the final set after $t$ steps. For simple settings like linear dynamical systems and convex polytopic initial sets reachability analysis is straightforward and fast [1]. However, many modern control systems operate on nonlinear systems [2], and increasingly contain neural network components. Reachability methods for such complex systems may not exist, and if they do they they almost always require under or over-approximation (also referred to as inner and outer approximation) of the reachable set as the exact set is not computable. It is desirable to produce tight reachable sets which have as little approximation error as possible, but this often trades off with tractability – the computation time needed to compute the reachable set.
@@
Figure 1: An illustration of overapproximate forward reachability. A symbolic query implicitly constrains the reachable set at intermediate timesteps producing tighter final sets. This concept is central to the refinement algorithm.
@@
There have been many approaches in the reachability literature to address tractability and tightness. One of the central concepts used is domain refinement [3, 4]. Domain refinement refers to splitting the domain into subsets, e.g., through gridding, and computing the reachable set for each subset. Without domain refinement, the reachable set approximation would be unusably loose, but if gridded too finely the problem can become intractable as there is then exponential complexity in the number of state dimensions.
@@
Another concept that has been used in the reachability literature is decomposition [5, 6, 7]. Decomposition refers to re-writing the system dynamics into a collection of uncoupled or loosely coupled subsystems. If e.g., using gridding to produce tight set approximations, fine gridding can then be used on each low dimensional subsystem without incurring the exponential cost of gridding the original high dimensional system.
@@
Lastly, many methods make use of pre-specified template sets in order to speed up computation [8, 9, 10, 11]. In these methods, the tightest circumscribing (or inscribing) set of a particular type such as a hyperrectangle, ellipsoid, or zonotope is computed rather than computing the arbitrary non-convex true reachable set.
@@
In this work we make use of a concept known as symbolic reachability [8] to define an algorithm for temporal refinement. Many verification methods use concrete reachability which involves computing the approximate reachable set at time $t+1$ using only the approximate reachable set at time $t$, or $t-1$ from $t$ in the backward case [12]. While concrete reachability is fast, it accumulates large approximation error quickly, a phenomenon known as the “wrapping effect” [13]. Symbolic reachability can be used to mitigate this effect [8]. Using symbolic reachability, the approximate reachable set at e.g., time $t+2$ is computed from the set at time $t$ while implicitly constraining the system trajectories to pass through the intermediate set at time $t+1$ (fig. 1). This results in tighter sets than if the set at $t+2$ was computed using only the approximate set at time $t+1$. It is possible to perform symbolic reachability for limited time horizons (temporal depths), each iteration computing a set farther into the future (or further in the past) and implicitly constraining system trajectories to pass through all sets at timesteps in between, but the problem will quickly approach intractability. To address this tradeoff between tightness and tractability, the authors of [8] perform hybrid-symbolic reachability which uses mixtures of concrete and symbolic computations in a hand-tuned, ad-hoc manner.
@@
In this paper, we contribute a temporal refinement algorithm to automate hybrid-symbolic reachability analysis for tenuously tractable discrete-time dynamical systems. The temporal depth and end timestep of each query are automatically selected to produce tight reachable sets given the remaining computational budget and projected total computation length. Our algorithm is an orthogonal contribution to existing strategies for balancing tightness and tractability because it may be combined with concepts such as domain refinement, system decomposition, and pre-specified template sets; among others.  [^1] Note that while temporal refinement may sound similar to timestep adaptation for continuous systems [14], it is not closely related. In a continuous time system, the length of time step $\Delta t$ is an algorithm hyperparameter that maybe adapted according to error requirements but in a discrete time system, there is either no specific step length, only time indices, $t=0$, $t=1$ or there is a fixed timestep length $\Delta t$ which cannot be adapted.
@@
We demonstrate our algorithm for temporal refinement on overapproximate forward reachability analysis, but it is similarly applicable to backward reachability analysis and underapproximation. Specifically, we treat a challenging class of problems: computing forward reachable sets for discrete time nonlinear dynamical systems with neural network control policies; also called Neural Feedback Loops (NFLs). For this class of problem, there has been work on computing forward reachable sets [8], backward reachable sets [15], performing domain refinement [16, 4, 17], and using exotic template polyhedron [18] but there remains difficulty balancing tractability and tightness making it a good class of problem on which to demonstrate our technique. We use several examples from the literature and show that our approach can produce a range of overapproximation error values depending on the allotted computational budget, including computing reachable sets 20-70% faster than previous hand-tuned approaches for the same amount of error in the reachable set approximation.
2 Problem Definition
@@
where $\mathcal{X}_{0}$ is the initial set of states. As aforementioned, methods exist to compute the exact reachable set $\mathcal{R}_{t}$ for classes of linear or piecewise linear systems, but for arbitrary smooth nonlinear dynamics, approximation (over or under) must be used. In this paper we denote the overapproximate reachable set $\tilde{\mathcal{R}_{t}}$ where $\mathcal{R}_{t}\subseteq\tilde{\mathcal{R}_{t}}$. The overapproximation error $e$ at a given time $t$ is defined
@@
Thus the problem this paper addresses may be stated as follows. When performing finite time hybrid-symbolic reachability analysis of $n$ steps for a discrete time dynamical system, algorithmically select points $t_{i}\in\{1,\ldots,n\}$ at which to perform symbolic reachability queries and select temporal depths $b_{\text{steps}}\in\{1\ldots t_{i}\}$ for those symbolic queries so as to minimize approximation error (eq. 3) while adhering to a pre-specified computational budget $b$.
Figure 2: Symbolic reachability scheme for primary refinement. In this example $b_{\text{steps}}=4$.
3 Refinement Algorithm
@@
At a high level, the algorithm begins in the “search” phase and searches for an estimate of the longest tractable temporal depth (symbolic horizon) $b_{\text{steps}}$. The algorithm then moves to a “jump” phase and performs symbolic queries of that temporal depth or shorter until the set at the desired final time $n$ has been reached. Alg. 1 provides more detail. In alg. 1, $\mathcal{X}_{0}$ is the initial set for reachability analysis, $b$ is the time budget in seconds, and $q$ is a query object consisting of a dynamics function $f$, a neural network controller $c(x):=NN(x)$, a final time horizon $n$, a variable to store current temporal depth $h$, and various parameters (e.g., hyperparameters controlling tightness of the nonlinear overapproximation). The method returns $\tilde{\mathbb{R}}$ a vector of reachable sets ~$\tilde{\mathcal{R}}_{t},\forall t\in[1\ldots n]$.
@@
The motivation for the structure of the algorithm comes partially from the initial work on hybrid-symbolic reachability [8] which used long symbolic jumps with one-step concrete queries to generate the sets in between. Long symbolic jumps produce tight approximations of the reachable set and beginning symbolic jumps from tight approximations limits the wrapping effect. However, in this paper, a reasonable symbolic “jump” size is not known a priori so we first search for a long jump size. The search procedure also has the benefit of limiting reachable set growth. Small errors early in the time horizon compound over many timesteps, and because the search phase produces tight reachable sets for early steps, it limits error compounding.
@@
Alg. 1 begins in the “search” phase with a temporal depth of $b_{\text{steps}}=1$ (line 3) beginning from the set $\mathcal{X}_{0}$ at time $t_{0}$ (lines 4-5). A symbolic reachability query is then performed (lines 7-8), and the new set(s) are then logged (lines 9-13). If there is enough time budget to continue searching, $b_{\text{steps}}$ will increment by $1$ (line 16). Symbolic reachable set computation queries are performed from the initial set $\mathcal{X}_{0}$ at time $t_{0}$ to increasingly large temporal depths / numbers of steps $b_{\text{steps}}$ into the future until it is estimated that trying a temporal depth one step longer would put the entire reachable set procedure over budget. Once this occurs, the algorithm will enter the “jump” phase and perform symbolic queries beginning from the set at $t_{\text{start}}:=t_{\text{cur}}$ (line 18-19) of temporal depth $b_{\text{steps}}$ or shorter until the set at the desired final time $n$ has been reached (line 6). Line 21 keeps $b_{\text{steps}}$ within correct range; and line 22 updates the early stopping timeout. The early stopping timeout limits budget overruns and is hand-tuned. The algorithm returns the reachable sets from time $t=1\ldots n$.
1 function RefinedReach($q$, $\mathcal{X}_{0}$, $b$):
@@ 25            $b_{\text{steps}}\leftarrow\min(b_{\text{steps}_{0}},n-t_{\text{start}})$
One of the key pieces of Algorithm 1 lies within the subroutine calc_steps (line 16) which determines whether to continue searching for a longer temporal depth $b_{\text{steps}}$ or to move to the “jump” phase. The complete algorithm for calc_steps is included in the Appendix (alg. 2) and described here. calc_steps first estimates the amount of time needed compute one symbolic step, $t_{\text{est}}$, assuming that the time to compute a symbolic query of $m$ steps is roughly $t_{\text{est}}*m$. This is a simplification that is not strictly accurate – the time a symbolic query takes may be in the worst case exponential in temporal depth for some problems and timing depends on the particular input set – but it leads to roughly accurate budgeting given that $t_{\text{est}}$ is updated at each iteration. The algorithm then estimates the time that future queries will take and uses this to determine if there is enough time budget to continue searching. This time estimate includes a symbolic query from $t_{0}$ to $b_{\text{steps}}+1$ plus the time to finish finish the temporal horizon in the “jump” phase. If the remaining budget $b$ is large enough, the algorithm increments $b_{\text{steps}}\leftarrow b_{\text{steps}}+1$ and if not, the algorithm enters the “jump” phase. Finally, if the solver stops early, the phase is also switched to “jump”. The early stopping is used to ensure the budget is approximately respected given that the estimated time per symbolic step $t_{\text{est}}$ is not an upper bound.
Note that long symbolic queries generally produce tight reachable sets but shorter queries can produce tighter sets if the starting set of the shorter query is tighter itself.
4 Overapproximate Forward Reachability
@@
Our refinement algorithm is applicable to both forwards and backwards reachability and both over (outer) and under (inner) approximation but in this paper, we demonstrate our refinement algorithm on overapproximate forward reachability of nonlinear neural feedback loops (NFLs). In particular, we use the reachability approach presented in [8]. To briefly summarize, the approach computes tight piecewise linear bounds for smooth nonlinear functions in the dynamics and then encodes these bounds as well as the ReLU neural network control policy into a mixed integer linear program. A hyperrectangular template set is then optimized to obtain an overapproximation of the reachable set. The mixture of symbolic and concrete queries is determined by the refinement algorithm presented here, but the procedure to approximate nonlinear functions and compute approximate reachable sets using mixed integer programs comes from [8] where further details can be found.
5 Soundness, Completeness, Complexity
The algorithm is sound, i.e. $\mathcal{R}_{t}\subseteq\tilde{\mathcal{R}}_{t}$, for all $t$, following directly from the soundness of [8]. In terms of completeness, the algorithm presented here does not guarantee that the overapproximate reachable sets ~$\tilde{\mathcal{R}}_{t}$ converge to the true reachable sets $\mathcal{R}_{t}$. The method presented here could in theory be combined with methods that provably converge to the true reachable set for some classes of nonlinear systems [19] to overcome this limitation, but completeness in this sense is beyond the scope of this paper.
The algorithm is guaranteed to produce a sequence of reachable sets as even with early stopping, the algorithm will extend the compute time until a finite feasible solution with $<50\%$ relative duality gap has been found. The algorithm is guaranteed to terminate under the mild assumption that a feasible solution with $<50\%$ relative duality gap exists.
The time complexity of this algorithm is in the worst case NP-hard as it involves solving mixed integer linear programs. However, the average case complexity is in practice more reasonable as is demonstrated in the results section.
6 Numerical Experiments
Numerical experiments were run on examples taken from  [8, 20] to assess the tightness and computational speed of the temporal refinement algorithm. Four problems were used: the pendulum dynamical system with a controller of 2 layers, 25 neurons per layer (S1); the TORA dynamical system with a controller of 3 layers, 25 neurons per layer (T1); the car dynamical system with a controller of 1 layer with 100 neurons (C1) and a controller of 1 layer with 200 neurons (C2). The reader is referred to [8] for the details on the neural network controlled dynamical systems.
6.1 Results
The temporal refinement algorithm was run multiple times to assess how changing the time budget changes the approximate reachable set error. Figure 3 measures the approximation error of the sets computed by the optimizer using the ratio of the sum of volumes:

### v2

@@ Abstract
Reachable set computation is an important tool for analyzing control systems. Simulating a control system can show general trends, but a formal tool like reachability analysis can provide guarantees of correctness. Reachability analysis for complex control systems, e.g., with nonlinear dynamics and/or a neural network controller, is often either slow or overly conservative. To address these challenges, much literature has focused on spatial refinement, i.e., tuning the discretization of the input sets and intermediate reachable sets. This paper introduces the idea of temporal refinement: automatically choosing when along the horizon of the reachability problem to execute slow symbolic queries which incur less approximation error versus fast concrete queries which incur more approximation error. Temporal refinement can be combined with other refinement approaches as an additional tool to trade off tractability and tightness in approximate reachable set computation. We introduce a temporal refinement algorithm and demonstrate its effectiveness at computing approximate reachable sets for nonlinear systems with neural network controllers. We calculate reachable sets with varying computational budget and show that our algorithm can generate approximate reachable sets with a similar amount of error to the baseline in 20-70% less time.
I Introduction
@@
Every controller needs to be analyzed for performance and reliability. Simulation is useful to get an idea of the general behavior of a control policy, but only formal analysis can guarantee correctness with respect to a model of the system and a specification. In this paper, we focus on reach-avoid specifications which dictate a goal set the system must reach and/or an avoid set it must not reach. For a given discrete time dynamical system with a control policy and set of possible starting states the forward reachable set at time $t$ is the set of states that the system could reach $t$ steps into the future. Given a final set instead of an initial set, the $t$-step backward reachable set is the set of states that will reach the final set after $t$ steps. For simple settings like linear dynamical systems and convex polytopic initial sets reachability analysis is straightforward and fast [frehse2011spaceex]. However, many modern control systems operate on nonlinear systems [slotine1991applied], and increasingly contain neural network components. Reachability methods for such complex systems may not exist, and if they do, they almost always require under or over-approximation (also referred to as inner and outer approximation) of the reachable set as the exact set is not computable. It is desirable to produce tight reachable sets which have as little approximation error as possible, but this often trades off with tractability – the computation time needed to compute the reachable set.
@@
Figure 1: An illustration of overapproximate forward reachability. A symbolic query uses multiple copies of the dynamical system update function $f(x)$ within a single computation to produce tighter final sets. This concept is central to the refinement algorithm.
@@
There have been many approaches in the reachability literature to address tractability and tightness. One of the central concepts used is domain refinement [bansal2017hamilton, everett2023drip]. Domain refinement refers to splitting the domain into subsets, e.g., through gridding, and computing the reachable set for each subset. Without domain refinement, the reachable set approximation would be unusably loose, but if gridded too finely the problem can become intractable as there is then exponential complexity in the number of state dimensions.
@@
Another concept that has been used in the reachability literature is decomposition [chen2018decomposition, mitchell2003overapproximating, chen2016decomposed]. Decomposition refers to re-writing the system dynamics as several uncoupled or loosely coupled subsystems. If using gridding to produce tight set approximations, fine gridding can then be used on each low dimensional subsystem without the exponential cost of gridding the original high dimensional system.
@@
Lastly, many methods make use of pre-specified template sets in order to speed up computation [sidrane2022overt, kurzhanski2000ellipsoidal, kuhn1998rigorously, althoff2011zonotope]. In these methods, the tightest circumscribing (or inscribing) set of a particular type such as a hyperrectangle, ellipsoid, or zonotope is computed rather than computing the arbitrary non-convex true reachable set.
@@
In this work we make use of a concept known as symbolic reachability [sidrane2022overt] to define an algorithm for temporal refinement. Many discrete-time verification methods use concrete reachability which involves computing the approximate reachable set at time $t+1$ using only the approximate reachable set at time $t$ (or $t-1$ from $t$ in the backward case) and a single copy of the dynamical system update function [siefert2023successor]. While concrete reachability is fast, it accumulates large approximation error quickly, a phenomenon known as the “wrapping effect” [neumaier1993wrapping]. Symbolic reachability can be used to mitigate this effect [sidrane2022overt]. Using symbolic reachability, the approximate reachable set at e.g., time $t+2$ is computed from the set at time $t$ using e.g., two copies of the dynamical system update function (see fig. 1). This results in tighter sets than if the set at $t+2$ was computed using only the approximate set at time $t+1$ and a single copy of the update function. It is possible to perform symbolic reachability for limited time horizons (temporal depths), each iteration computing a set farther into the future (or further in the past), but the problem will quickly approach intractability. To address this tradeoff between tightness and tractability, the authors of [sidrane2022overt] perform hybrid-symbolic reachability which uses mixtures of concrete and symbolic computations in a hand-tuned, ad-hoc manner.
@@
In this paper, we contribute a temporal refinement algorithm to automate hybrid-symbolic reachability analysis for tenuously tractable discrete-time dynamical systems. The temporal depth and end timestep of each query are automatically selected to produce tight reachable sets given the approximate remaining computational budget and approximate projected total computation length. Our algorithm is an orthogonal contribution to existing strategies for balancing tightness and tractability because it may be combined with concepts such as domain refinement, system decomposition, and pre-specified template sets; among others.  [^1] Note that while temporal refinement may sound similar to timestep adaptation for continuous systems [wetzlinger2022adaptive], it is not closely related. In a continuous time system, the length of time step $\Delta t$ is an algorithm hyperparameter that maybe adapted according to error requirements, but in a discrete time system, there is either no specific step length, only time indices, $t=0,t=1,\ldots$ or there is a fixed timestep length $\Delta t$.
@@
We demonstrate our algorithm for temporal refinement on overapproximate forward reachability analysis, but it is similarly applicable to backward reachability analysis and underapproximation. Specifically, we treat a challenging class of problems: computing forward reachable sets for discrete time nonlinear dynamical systems with neural network control policies; also called Neural Feedback Loops (NFLs). For this class of problem, there has been work on computing forward reachable sets [sidrane2022overt], backward reachable sets [rober2023backward], performing domain refinement [rober2023hybrid, everett2023drip, xiang2020reachable], and using exotic template polyhedron [manzanas2021verification] but there remains difficulty balancing tractability and tightness making it a good class of problem on which to demonstrate our technique. We use several examples from the literature and show that our approach can produce a range of overapproximation error values depending on the allotted approximate computational budget, including computing reachable sets 20-70% faster than previous hand-tuned approaches for the same amount of error in the reachable set approximation.
II Problem Definition
@@
[…]
The overapproximation error $e$ at a given time $t$ is defined
@@
Thus, the problem may be stated as follows.
Problem 1: Compute a vector of $n$ approximate reachable sets ~[~$\tilde{\mathbb{R}}~{}\triangleq~{}[\tilde{\mathcal{R}}_{1},\ldots,\tilde{\mathcal{R}}_{n}]$ using hybrid-symbolic reachability analysis with an automatically chosen mixture of concrete and symbolic queries so as to minimize approximation error (eq. 3) while approximately adhering to a pre-specified computational budget $b$.
III Refinement Algorithm
@@
We introduce Alg. 1, illustrated in fig. 2, as a solution to Problem 1. Alg. 1 automates hybrid symbolic reachability analysis by automatically selecting points $t_{i}\in\{1,\ldots,n\}$ at which to perform symbolic reachability queries, and temporal depths $b_{\text{steps}}\in\{1\ldots t_{i}\}$ for those symbolic queries so as to minimize approximation error. At a high level, the algorithm begins in the “search” phase and searches for an estimate of the longest tractable temporal depth (symbolic horizon) $b_{\text{steps}}$. The algorithm then moves to a “jump” phase and performs symbolic queries of that temporal depth or shorter until the set at the desired final time $n$ has been reached. More detail is provided in Alg. 1, where $\mathcal{X}_{0}$ is the initial set for reachability analysis, $b$ is the approximate time budget in seconds, symbolic_reach is a chosen symbolic reachability procedure, and $q$ is a query object consisting of the transition system and reachability horizon. For our chosen class of example problems, $q$ consists of a nonlinear dynamics function $f$, a neural network controller $c(x):=NN(x)$, a final time horizon $n$, a variable to store current temporal depth $h$, and hyperparameters, e.g., controlling tightness of the nonlinear overapproximation. The method returns $\tilde{\mathbb{R}}$ a vector of approximate reachable sets ~$\tilde{\mathcal{R}}_{t},\forall t\in[1\ldots n]$.
Figure 2: An illustration of Alg. 1. In this example $b_{\text{steps}}=4$.
@@
The motivation for the structure of the algorithm comes partially from the initial work on hybrid-symbolic reachability [sidrane2022overt] which used long symbolic jumps with one-step concrete queries to generate the sets in between. Long symbolic jumps produce tight approximations of the reachable set and beginning symbolic jumps from tight approximations limits the wrapping effect. However, in this paper, a reasonable symbolic “jump” size is not known a priori so we first search for a long jump size. The search procedure also has the benefit of limiting reachable set growth. Small errors early in the time horizon compound over many timesteps, and because the search phase produces tight reachable sets for early steps, it limits error compounding.
@@
Alg. 1 begins in the “search” phase with a temporal depth of $b_{\text{steps}}=1$ (line 3) beginning from the set $\mathcal{X}_{0}$ at time $t_{0}$ (lines 4-5). A symbolic reachability query is then performed (lines 7-8), and the new set(s) are then logged (lines 9-13). If there is enough time budget to continue searching, $b_{\text{steps}}$ will increment by $1$ (line 16). Symbolic reachable set computation queries are performed from the initial set $\mathcal{X}_{0}$ at time $t_{0}$ to increasingly large temporal depths / numbers of steps $b_{\text{steps}}$ into the future until it is estimated that trying a temporal depth one step longer would put the entire reachable set procedure over budget. Once this occurs, the algorithm will enter the “jump” phase and perform symbolic queries beginning from the set at $t_{\text{start}}:=t_{\text{cur}}$ (line 18-19) of temporal depth $b_{\text{steps}}$ or shorter until the set at the desired final time $n$ has been reached (line 6). Line 21 keeps $b_{\text{steps}}$ within correct range; and line 22 updates the early stopping timeout. The early stopping timeout limits budget overruns and is hand-tuned. The algorithm returns reachable sets from $t=1\ldots n$.
1 function RefinedReach($q$, $\mathcal{X}_{0}$, $b$, symbolic_reach):
@@ 25            $b_{\text{steps}}\leftarrow\min(b_{\text{steps}_{0}},n-t_{\text{start}})$
One of the key pieces of Alg. 1 is the subroutine calc_steps (line 16) which determines whether to continue searching for a longer temporal depth $b_{\text{steps}}$ or to move to the “jump” phase. The complete algorithm for calc_steps is included in the Appendix (alg. LABEL:alg:calc_qsteps) and described here. calc_steps first estimates the time needed to compute one symbolic step, $t_{\text{est}}$, assuming that the time to compute a symbolic query of $m$ steps is roughly $t_{\text{est}}*m$. This is a simplification that is not strictly accurate – the time a symbolic query takes may be in the worst case exponential in temporal depth, and timing depends on the particular input set – but it leads to roughly accurate budgeting when $t_{\text{est}}$ is updated at each iteration. The algorithm then estimates the time that future queries will take and uses this to determine if there is enough time budget to continue searching. This time estimate includes a symbolic query from $t_{0}$ to $b_{\text{steps}}+1$ plus the time to finish finish the temporal horizon in the “jump” phase. If the remaining budget $b$ is estimated to be large enough, the algorithm increments $b_{\text{steps}}\leftarrow b_{\text{steps}}+1$ and if not, the algorithm enters the “jump” phase. If the solver stops early, the phase is also switched to “jump”. Early stopping is used to ensure the budget is approximately respected given that the estimated time per symbolic step $t_{\text{est}}$ is not an upper bound. Note that long symbolic queries generally produce tight reachable sets but shorter queries can produce tighter sets if the starting set of the shorter query is tighter itself.
IV Overapproximate Forward Reachability
@@
Our refinement algorithm is applicable to both forwards and backwards reachability and both over (outer) and under (inner) approximation, but in this paper we demonstrate our refinement algorithm on overapproximate forward reachability of nonlinear neural feedback loops (NFLs). In particular, we use the reachability approach presented in [sidrane2022overt]. To briefly summarize, the approach rewrites high dimensional nonlinear dynamics into many one dimensional functions. It then computes tight piecewise linear bounds for each smooth nonlinear 1d function and then encodes these bounds as constraints in a mixed integer linear program (MILP). The neural network control policy is limited to ReLU activation functions making it a piecewise linear function that is then encoded as constraints in the MILP as well. A hyperrectangular template set is then optimized to obtain an overapproximation of the reachable set. The hyperrectangular set is constructed by finding the minimum and maximum value of each state variable $\vec{x}^{(i)}_{t},i\in 1\ldots d$. The mixture of symbolic and concrete queries is determined by the refinement algorithm presented here, but the procedure to approximate nonlinear functions and compute approximate reachable sets using MILPs comes from [sidrane2022overt] where further details can be found.
V Soundness, Completeness, Complexity
The section serves to demonstrate the soundness, completeness in the sense of termination, and complexity of the proposed Algorithm 1 for solving Problem 1. We first show two supporting lemmas, V.1 and V.2, before asserting the main theorem V.3.
Lemma V.1
Alg. 1 produces sets ~$\tilde{\mathcal{R}}_{t}\in\tilde{\mathbb{R}}$ which overapproximate the true reachable set $\mathcal{R}_{t}\subseteq\tilde{\mathcal{R}}_{t}$ at each time $t\in\{1,\ldots,n\}$.
Proof:
This follows directly from the overapproximation property of the symbolic_reach procedure being called. E.g., in the examples presented here, from [sidrane2022overt]. $\square$